OpenStudy (anonymous):

an object has a mass of 100 kg and had an initial velocity of 3 m/s then 50 kg was added to this object. what's the new velocity of the item?

OpenStudy (anonymous):

What is the distance that the object travels?

OpenStudy (anonymous):

To reply this question in a decent way, we need much more details. We have acceleration? We have the initial conditions about the second object like direction and initial velocity?

OpenStudy (anonymous):

well to be more specific then. the question reads: an ice skater with a mass of 100kg is skating past his younger brother who is at rest on the ice. As the ice skater passes his younger brother, he picks his younger brother off the ground. If his younger brother has a mass of 50 kg and the ice skater initially has a velocity of 3 m/s, then what must be their new velocity as they skate away together?

OpenStudy (anonymous):

2m/s

OpenStudy (anonymous):

U1=3m/s , M1=100Kg, V=? U2=0m/s, M2=50Kg consider this as a perfect inelastic collision.....then M1U1+M2U2=(M1+M2)V V=300/150=2m/s